# VOCAL: Visual Odometry via ContrAstive Learning

### v1

@@ 1 Introduction
Figure 1: (a) Conventional graph-based visual odometry, where connections between camera states $x_{i}$ and features $f_{j}$ are modeled using predefined graphs $z_{ij}$. This manual design limits both flexibility and interpretability in learning-based VO systems. (b) The VOCAL architecture eliminates the need for handcrafted graph structures by reframing VO as a label-ranking problem. Through contrastive learning, VOCAL organizes features extracted from visual inputs based on their corresponding camera states, ensuring that inputs with similar camera states yield consistent features in the latent space. This approach improves spatial understanding in visual odometry by establishing a direct correlation between feature representations and 3D camera states.
@@ 3.3 Label Ranking in Learning-based VO
Thus, the ranking-based probability model is directly linked to learning-based VO, enabling VOCAL to organize observations by their camera states.
@@ 3.4 Visual Odometry via Contrastive Learning
$l^{i}_{SupRNC}={}$ $\frac{1}{2N-1}\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{2N}\Biggl{[}-\log\Bigl{(}\frac{\exp\Bigl{(}sim(f_{z_{i}},f_{z_{j}})/\tau\Bigr{)}}{\sum_{f_{z_{k}}\in S_{i,j}}\exp\Bigl{(}sim(f_{z_{i}},f_{z_{k}})/\tau\Bigr{)}}\Bigr{)}\Biggr{]}$ (10)
$\quad+\lambda\,L1\bigl{(}x_{i},\hat{x}_{i}\bigr{)}.$
@@ 4 Experiments
[…]
@@
Dimension $roll$ $pitch$ $yaw$
@@
Table 1: Correlations between Features and Camera States: High correlation scores for $z$ translation and $pitch$ rotation reflect the dominant motion in the KITTI dataset. In contrast, the $y$ and $yaw$ dimensions—typically regarded as noise in this dataset—exhibit lower correlation scores for both $r_{s}$ and $r_{k}$.
@@ 4.1 Feature Ranking
We analyze the feature distribution produced by our contrastive feature encoder in the latent space. By leveraging the Supervised Rank-N-Contrast loss ($L_{SupRNC}$), the learned features are expected to closely correlate with the ground-truth camera states. Table 1 reports the Spearman ($r_{s}$) and Kendall ($r_{k}$) rank correlation coefficients, where values approaching 1.0 indicate more effective ranking. Notably, the $z$-translation and $pitch$-rotation achieve high correlation scores, reflecting the dominant motions along the $z$-axis and changes in $pitch$ in the KITTI dataset. In contrast, the $y$ and $yaw$ dimensions, generally considered noise, exhibit lower correlation scores.
Fig. 5 illustrates the feature distribution for $z$ and $pitch$ (distributions for the remaining dimensions are provided in the supplementary material Sec. 8). In the figure, lighter colors (yellow) represent larger camera state values, while darker hues (purple) indicate smaller values. The continuous gradient from light to dark underscores the effective ranking of features by camera state, yielding an interpretable and flexible latent representation that also supports the estimation of the most likely camera states.
@@ 4.2 Generalization Capability
Method Training dataset KITTI Sequences ($\mathrm{t}_{rel}$/$\mathrm{r}_{rel}$)
@@
VOCAL (ours) KITTI 00, 02, 08, 09 4.76/1.98 5.03/1.77 5.67/2.45 24.32/9.26 12.61/4.82 14.80/8.13
Table 2: Visual Odometry Results ($\mathrm{t}_{rel}$/$\mathrm{r}_{rel}$) on KITTI Sequences: $\mathrm{t}_{rel}$ denotes the average translational RMSE drift (in %) over trajectory segments of 100–800 m, and $\mathrm{r}_{rel}$ denotes the average rotational RMSE drift (in /100 m) over trajectory segments of 100–800 m. An asterisk () marks training data (which may be overfitted).
@@
Fig. 6 plots the rank correlation for all six degrees of freedom (DoF) as a function of the training data proportion. Notably, the $z$ and $pitch$ dimensions—representing the dominant motions in the KITTI dataset—exhibit gradually increasing $r_{s}$ values as more training data becomes available, indicating that our method ranks features more effectively with additional data. In contrast, the $y$ and $yaw$ dimensions, which are more susceptible to noise in KITTI, do not converge regardless of the training set size. Additionally, Fig. 6 presents the performance of $\mathrm{t}_{rel}$ and $\mathrm{r}_{rel}$ on different test datasets, showing that our model performs better as the training data scale increases. Even with a relatively small training set, VOCAL achieves competitive performance compared with other methods (as will be detailed in the next section). This early-stage performance is due to our method’s ability to directly capture the relationship between visual features and camera states through label ranking.
@@ 4.3 VO Performance
Our experimental results show that VOCAL achieves the best translation performance on sequence 03 and ranks second on sequence 05. For rotation accuracy, VOCAL attains the best performance on sequences 03, 04, and 05. Notably, although our training data is relatively smaller than that used in [45, 42], VOCAL still performs competitively — even outperforming those methods on training sequences 03–05 (which may be overfitted). These findings demonstrate that VOCAL not only provides an interpretable latent representation but also delivers competitive VO performance.
@@ 5 Conclusion
6 Acknowledgment
We thank Peng Cao (CSAIL, MIT) for detailed explanations of Rank-N-Contrast and valuable suggestions, and Sangmin Jung (SCAI, ASU) for assistance with experiments. We also gratefully acknowledge the support and resources provided by the Research Computing facilities at Arizona State University [12]. This work is supported by the Toyota Research Institute of North America (Mothership Project), and Chi-Yao Huang is additionally supported by the ASU Fulton PhD Fellowship.
@@ 7 Resources
For additional resources and demo videos, please visit: https://anonymous.4open.science/r/vocal-3315/.
@@ 8 Feature Distribution
9 Trajectory
In this section, we present the 3D trajectories for the test datasets (KITTI sequences 03–07 and 10).
Figure 13: KITTI 03 Trajectory
Figure 14: KITTI 04 Trajectory
Figure 15: KITTI 05 Trajectory
Figure 16: KITTI 06 Trajectory
Figure 17: KITTI 07 Trajectory
Figure 18: KITTI 10 Trajectory

### v2

@@ 1 Introduction
Figure 1: (a) Conventional graph-based visual odometry, where connections between camera states $x_{i}$ and features $f_{j}$ are modeled using predefined graphs $z_{ij}$. This manual design limits both flexibility and interpretability in learning-based VO systems. (b) The VOCAL architecture eliminates the need for handcrafted graph structures by reframing VO as a label-ranking problem. Through contrastive learning, VOCAL organizes features extracted from visual inputs based on their corresponding camera states, ensuring that inputs with similar camera states yield consistent features in the latent space. Our approach improves spatial understanding in visual odometry by establishing a direct correlation between feature representations and 3D camera states.
@@ 3.3 Label Ranking in Learning-based VO
The ranking-based probability model is linked to learning-based VO, enabling VOCAL to organize observations by their camera states.
@@ 3.4 Visual Odometry via Contrastive Learning
$l^{i}_{SupRNC}={}$ (10)
$\frac{1}{2N-1}\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{2N}\Biggl[-\log\Bigl(\frac{\exp\Bigl(sim(f_{z_{i}},f_{z_{j}})/\tau\Bigr)}{\sum_{f_{z_{k}}\in S_{i,j}}\exp\Bigl(sim(f_{z_{i}},f_{z_{k}})/\tau\Bigr)}\Bigr)\Biggr]$ (11)
$\quad+\lambda\,L1\bigl(x_{i},\hat{x}_{i}\bigr).$ (12)
@@ 4 Experiments
Prior to training, we compute the optical flow for each pair of consecutive images using the Gunnar-Farneback method. We set the parameters in the OpenCV function as follows: $\text{pyr\_scale}=0.5$, $\text{levels}=3$, $\text{winsize}=15$, $\text{poly\_n}=5$, and $\text{poly\_sigma}=1.2$. The resulting flow maps are then cropped to $224\times 224$. We apply Gaussian noise augmentation with mean $\mu=0.0$ and standard deviation $\sigma=0.05$.
@@
Dimension rollroll pitchpitch yawyaw
@@
Table 1: Correlations between Features and Camera States: High correlation scores for $z$ translation and pitchpitch rotation reflect the dominant motion in the KITTI dataset. In contrast, the $y$ and yawyaw dimensions—typically regarded as noise in this dataset—exhibit lower correlation scores for both $r_{s}$ and $r_{k}$.
@@ 4.1 Feature Ranking
We analyze the feature distribution produced by our contrastive feature encoder in the latent space. By leveraging the Supervised Rank-N-Contrast loss ($L_{SupRNC}$), the learned features are expected to closely correlate with the ground-truth camera states. Table 1 reports the Spearman ($r_{s}$) and Kendall ($r_{k}$) rank correlation coefficients, where values approaching 1.0 indicate more effective ranking. Notably, the $z$-translation and pitchpitch-rotation achieve high correlation scores, reflecting the dominant motions along the $z$-axis and changes in pitchpitch in the KITTI dataset. In contrast, the $y$ and yawyaw dimensions, generally considered noise, exhibit lower correlation scores.
Fig. 5 illustrates the feature distribution for $z$ and pitchpitch (distributions for the remaining dimensions are provided in the supplementary material Sec.˜7). In the figure, lighter colors (yellow) represent larger camera state values, while darker hues (purple) indicate smaller values. The continuous gradient from light to dark underscores the effective ranking of features by camera state, yielding an interpretable and flexible latent representation that also supports the estimation of the most likely camera states.
@@ 4.2 Generalization Capability
Method Training dataset KITTI Sequences ($\mathrm{t}_{rel}$$\mathrm{r}_{rel}$)
@@
VOCAL (ours) KITTI 00, 02, 08, 09 4.60/1.99 5.18/1.51 5.45/2.51 5.11/6.06 12.13/4.79 7.72/7.61
Table 2: Visual Odometry Results ($\mathrm{t}_{rel}$$\mathrm{r}_{rel}$) on KITTI Sequences: $\mathrm{t}_{rel}$ denotes the average translational RMSE drift (in %) over trajectory segments of 100–800 m, and $\mathrm{r}_{rel}$ denotes the average rotational RMSE drift (in /100 m) over trajectory segments of 100–800 m. An asterisk () marks training data (which may be overfitted).
@@
Fig. 6 plots the rank correlation for all six degrees of freedom (DoF) as a function of the training data proportion. Notably, the $z$ and pitchpitch dimensions—representing the dominant motions in the KITTI dataset—exhibit gradually increasing $r_{s}$ values as more training data becomes available, indicating that our method ranks features more effectively with additional data. In contrast, the $y$ and yawyaw dimensions, which are more susceptible to noise in KITTI, do not converge regardless of the training set size. Additionally, Fig. 6 presents the performance of $\mathrm{t}_{rel}$ and $\mathrm{r}_{rel}$ on different test datasets, showing that our model performs better as the training data scale increases. Even with a relatively small training set, VOCAL achieves competitive performance compared with other methods (as will be detailed in the next section). This early-stage performance is due to our method’s ability to directly capture the relationship between visual features and camera states through label ranking.
@@ 4.3 VO Performance
Our experimental results show that VOCAL achieves the best translation performance on sequence 03 and ranks second on sequences 05, 06, and 10. For rotation accuracy, VOCAL attains the best performance on sequences 03, 04, and 05. Notably, although our training data is relatively smaller than that used in [44, 41], VOCAL still performs competitively, even outperforming those methods on training sequences 03–05. These findings demonstrate that VOCAL not only provides an interpretable latent representation but also delivers competitive VO performance.
@@ 6 Resources
For additional resources and demo videos, please visit: https://github.com/huang-chiyao/vocal.